ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 219, 2017

Balogh, J. and A. Z. Wagner On the number of union-free families  437

Bavula, V. V. and T. Lu The quantum euclidean algebra and its prime spectrum

929

Bell, J. P., C. Ingalls and R. Ramkumar Embedding of quotient division algebras of rings of differential operators

441

Bennett, P. Normal forms for semigroup amalgams  379

Bridson, M. R. and D. H. Kochloukova The torsion-free rank of homology in towers of soluble pro-p groups

817

Burton, P. Naive entropy of dynamical systems  637

Butterley, O. and I. Melbourne Disintegration of invariant measures for hyperbolic skew products

171

Cavalcanti, M. M, V. N. Domingos Cavalcanti, M. A. Jorge Silva and C. M. Webler Exponential stability for the wave equation with degenerate nonlocal weak damping

189

Choi, S. and H. Park Wedge operations and a new family of projective toric manifolds

353

Chuaqui, M. Injectivity of minimal immersions and homeomorphic extensions to space

983

Coulembier, K. and V. Mazorchuk Dualities and derived equivalences for category O

661

Derksen, H., A. Van den Essen and W. Zhao The Gaussian Moments Conjecture and the Jacobian Conjecture

917

Domingos Cavalcanti, V. N. See M. M. Cavalcanti Esposito, C. and E. Miranda Rigidity of infinitesimal momentum

maps

757

Ferber, A., C. Lee and F. Mousset Packing spanning graphs from separable families

959

Glasner, Y. Invariant random subgroups of linear groups  215

Hackney, P. and M. Robertson The homotopy theory of simplicial props

835

Hänninen, T. S. Two-weight inequality for operator-valued positive dyadic operators by parallel stopping cubes

71

Hazrat, R., N. Vavilov and Z. Zhang Multiple commutator formulas for unitary groups

287

Ingalls, C. See J. P. Bell Jorge Silva, M. A. See M. M. Cavalcanti Kochloukova, D. H. See M. R. Bridson

Kopecká, E. and A. Paskiewicz Strange products of projections  271

Kramer, L. and J. Schillewaert Strongly transitive actions on euclidean buildings

163

Lascar, B. and R. Lascar Corrections to “L’inégalité de Melin-Hormander en charactéristiques multiples”, Israel Journal of Mathematics 211 (2016), 13–22

1013

Lascar, R. See B. Lascar Lee, C. See A. Ferber Lu, T. See V. V. Bavula Lü, J.-F., X.-F. Mao and J. J. Zhang The Nakayama automor-

phism of a class of graded algebras

707

Lucchini, A. and G. Tracey An upper bound on the Chebotarev invariant of a finite group

449

Mao, X.-F. See J.-F. Lü Mazorchuk, V. See K. Coulembier

Meiri, C. On the non-amenability of the reflective quotient  903

Melbourne, I. See O. Butterley Miranda, E. See C. Esposito G. Molica Bisci and V. D. Rădulescu A sharp eigenvalue theorem

for fractional elliptic equations

331

Mordhorst, O. New results on affine invariant points  529
Morra, S. Iwasawa modules and p-modular representations of GL2  1

Mousset, F. See A. Ferber

Naor, A. and Y. Rabani On Lipschitz extension from finite subsets  115

Nikolidakis, E. N. The Bellman function of the dyadic maximal operator related to Kolmogorov’s inequality

507

Park, H. See S. Choi Paskiewicz, A. See E. Kopecká Ponce, R. On the well-posedness of degenerate fractional differential

equations in vector valued function spaces

727

Pournin, L. The asymptotic diameter of cyclohedra  609

Rabani, Y. See A. Naor Rădulescu, V. D. See G. Molica Bisci Ramkumar, R. See J. P. Bell Robertson, M. See P. Hackney Schillewaert, J. See L. Kramer Shi, R. and B. Weiss Invariant measures for solvable groups and

Diophantine approximation

479

Smoktunowicz, A. How far can we go with Amitsur’s theorem in differential polynomial rings?

555

Snopce, I. and S. Tanushevski Test elements in pro-p groups with applications in discrete group

783

Tanushevski, S. See I. Snopce

Thom, A. About the length of laws for finite groups  469

Tracey, G. See A. Lucchini Van den Essen, A. See H. Derksen Vavilov, N. See R. Hazrat Wagner, A. Z. See J. Balogh Webler, C. M. See M. M. Cavalcanti Weiss, B. See R. Shi

White, M. J. On Tverberg partitions  549

Zhang, J. J. See J. F. Lü Zhang, Z. See R. Hazrat Zhao, W. See H. Derksen